# Observation of $D^{+}\to\bar{K}_{1}(1270)^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K_{1}(1270)^{-}\mu^{+}\nu_{\mu}$

### v1

@@ Abstract
Combining the BFs measured in this work and the BFs of ${\mathcal{B}}[D^{+}\to\bar{K}_{1}^{0}e^{+}\nu_{e}]=(2.30\pm{0.26\begin{subarray}{c}+0.18\\ -0.21\end{subarray}}\pm 0.25)\times 10^{-3}$ BESIII:2019dptok1 and ${\mathcal{B}}[D^{0}\to K_{1}^{-}e^{+}\nu_{e}]=(1.09\pm{0.13\begin{subarray}{c}+0.09\\ -0.13\end{subarray}}\pm 0.12)\times 10^{-3}$ BESIII:2021dotok1 reported in the previous BESIII analysis, we determine the BF ratios ($\mathcal{R}^{D}_{\mu/e}=\frac{{\mathcal{B}}[D\to\bar{K}_{1}\mu^{+}\nu_{\mu}]}{{\mathcal{B}}[D\to\bar{K}_{1}e^{+}\nu_{e}]}$) to be $\mathcal{R}^{D^{+}}_{\mu/e}=1.03\pm 0.14\begin{subarray}{c}+0.11\\ -0.15\end{subarray}$ and $\mathcal{R}^{D^{0}}_{\mu/e}=0.74\pm 0.13\begin{subarray}{c}+0.08\\ -0.13\end{subarray}$. Here, the systematic uncertainties in ST yields, $K^{-}/\pi^{\pm}$ tracking and PID, and $\pi^{0}$ reconstruction cancel. These are consistent with $\mu$-$e$ lepton flavor universality in $D^{+(0)}\to\bar{K}_{1}\ell^{+}\nu_{\ell}$ decays within current uncertainties. Using the BFs of $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ and $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ measured in this work as well as the world-average lifetimes of the $D^{0}$ and $D^{+}$ mesons pdg2022 , we determine the partial decay width ratio $\Gamma[D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}]/\Gamma[D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}]=1.22\pm 0.10\begin{subarray}{c}+0.06\\ -0.09\end{subarray}$, which is consistent with unity as predicted by isospin conservation.
I Acknowledgement
@@
The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2023YFA1606000, 2020YFA0406400, 2020YFA0406300; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11735014, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017, 12361141819; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; the CAS Center for Excellence in Particle Physics (CCEPP); Joint Large-Scale Scientific Facility Funds of the NSFC and CAS under Contract No. U1832207; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; German Research Foundation DFG under Contracts Nos. 455635585, FOR5327, GRK 2149; Istituto Nazionale di Fisica Nucleare, Italy; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B16F640076; Polish National Science Centre under Contract No. 2019/35/O/ST2/02907; The Swedish Research Council; U. S. Department of Energy under Contract No. DE-FG02-05ER41374.
@@
Figure 3 shows the distributions of $M_{K^{-}\pi^{+}\pi^{0}}$ from $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and $M_{K^{-}\pi^{+}\pi^{-}}$ for $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$ candidates, respectively. Results from the data and the inclusive MC sample are overlaid.
Fig. 3: Comparison of $M_{K^{-}\pi^{+}\pi^{0}}$ (left) and $M_{K^{-}\pi^{+}\pi^{-}}$ (right) of the candidates for $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$. The points with error bars are data, the blue filled histograms are the simulated background, and the red line histograms are the signal MC samples. A requirement of $|U_{\rm miss}|<0.02$ GeV has been applied.

### v2

@@ Abstract
Combining the BFs measured in this work and the BFs of ${\mathcal{B}}[D^{+}\to\bar{K}_{1}^{0}e^{+}\nu_{e}]=(2.30\pm{0.26\begin{subarray}{c}+0.18\\ -0.21\end{subarray}}\pm 0.25)\times 10^{-3}$ BESIII:2019dptok1 and ${\mathcal{B}}[D^{0}\to K_{1}^{-}e^{+}\nu_{e}]=(1.09\pm{0.13\begin{subarray}{c}+0.09\\ -0.13\end{subarray}}\pm 0.12)\times 10^{-3}$ BESIII:2021dotok1 reported in the previous BESIII analysis, we determine the BF ratios ($\mathcal{R}^{D}_{\mu/e}=\frac{{\mathcal{B}}[D\to\bar{K}_{1}\mu^{+}\nu_{\mu}]}{{\mathcal{B}}[D\to\bar{K}_{1}e^{+}\nu_{e}]}$) to be $\mathcal{R}^{D^{+}}_{\mu/e}=1.03\pm 0.14\begin{subarray}{c}+0.11\\ -0.15\end{subarray}$ and $\mathcal{R}^{D^{0}}_{\mu/e}=0.74\pm 0.13\begin{subarray}{c}+0.08\\ -0.13\end{subarray}$. Here, the systematic uncertainties in ST yields, $K^{-}/\pi^{\pm}$ tracking and PID, and $\pi^{0}$ reconstruction cancel. These are consistent with theoretical predictions based on $\mu$-$e$ LFU in $D^{+(0)}\to\bar{K}_{1}\ell^{+}\nu_{\ell}$ decays within current uncertainties. Using the BFs of $D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}$ and $D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}$ measured in this work as well as the world-average lifetimes of the $D^{0}$ and $D^{+}$ mesons pdg2022 , we determine the partial decay width ratio $\Gamma[D^{+}\to\bar{K}_{1}^{0}\mu^{+}\nu_{\mu}]/\Gamma[D^{0}\to K_{1}^{-}\mu^{+}\nu_{\mu}]=1.22\pm 0.10\begin{subarray}{c}+0.06\\ -0.09\end{subarray}$, which is consistent with unity as predicted by isospin conservation.
I Acknowledgements
@@
The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2023YFA1606000; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11735014, 11935015, 11935016, 11935018, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017, 12361141819; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; the CAS Center for Excellence in Particle Physics (CCEPP); Joint Large-Scale Scientific Facility Funds of the NSFC and CAS under Contract No. U1832207; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; German Research Foundation DFG under Contracts Nos. 455635585, FOR5327, GRK 2149; Istituto Nazionale di Fisica Nucleare, Italy; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B16F640076; Polish National Science Centre under Contract No. 2019/35/O/ST2/02907; The Swedish Research Council; U. S. Department of Energy under Contract No. DE-FG02-05ER41374.
@@
Figure 4 and 5 shows the distributions of $M_{K^{-}\pi^{+}}$, $M_{K^{-}\pi^{0(-)}}$ and $M_{\pi^{+}\pi^{0(-)}}$ for $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$ and $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$ candidates, respectively. Results from the data and the inclusive MC sample are overlaid.
Fig. 4: Comparison of $M_{K^{-}\pi^{+}}$ (left), $M_{K^{-}\pi^{0}}$ (center) and $M_{\pi^{+}\pi^{0}}$ (right) of the candidates for $D^{+}\to K^{-}\pi^{+}\pi^{0}\mu^{+}\nu_{\mu}$. The points with error bars are data, the blue filled histograms are the simulated background, and the red line histograms are the signal MC samples. A requirement of $|U_{\rm miss}|<0.02$ GeV has been applied.
Fig. 5: Comparison of $M_{K^{-}\pi^{+}}$ (left), $M_{K^{-}\pi^{-}}$ (center) and $M_{\pi^{+}\pi^{-}}$ (right) of the candidates for $D^{0}\to K^{-}\pi^{+}\pi^{-}\mu^{+}\nu_{\mu}$. The points with error bars are data, the blue filled histograms are the simulated background, and the red line histograms are the signal MC samples. A requirement of $|U_{\rm miss}|<0.02$ GeV has been applied.